Opened 7 years ago

Last modified 5 years ago

#341 assigned flight processing

EU10/01, 168/2010, Loobos

Reported by: jb362 Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: knpa

Description (last modified by emca)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EU10_01-2010_168_loobos

Data arrived from ARSF via network transfer on 23/06/2010

Scientific objective:

1.Key impact factors on CO2 and H2O fluxes
1.1 examine the influence of within-site elevation changes using flux tower observations, airborne LiDAR data, and ecosystem models data (Loobos)
1.2 develop an approach to allocate CO2 and H2O fluxes to specific tree species of a mixed forest (Norunda)
1.3 assess the influence of forest thinning on CO2 uptake (Norunda)
1.4 evaluate above findings and test transferability of findings (Tharandt)
1.5 investigate the impact of storm damage (Tharandt)

  1. Reducing uncertainty in the upscaling of greenhouse gases

2.1. develop an approach to evaluate airborne flux measurements using flux tower data, footprint model and airborne LiDAR data (Loobos)
2.2 use airborne LiDAR data to validate GLAS/ICESat for different forest biomes, where accurate tree height and vegetation cover density will be co-located with GLAS footprints
2.3 evaluate the ability of GLAS spatial sampling density to represent the wider field site
2.4 develop an approach to model CO2 and H2O fluxes on a regional scale using GLAS/ICESat data and generate key driving parameters for land-surface models
2.5 evaluate whether regional CO2/H2O flux models driven by satellite optical data can be improved by the addition of tree height information from satellite LiDAR

Priority: a5

PI: N. Kljun

Sensors:

  • Eagle (01/09/2010)
  • Hawk (21/12/2010)
  • Leica LIDAR (06/09/10)
  • RCD (26/08/10)

Change History (59)

comment:1 Changed 7 years ago by jb362

This directory needs to be backed up in the repository

comment:2 Changed 7 years ago by jb362

IPAS navigation for this project has been processed.

comment:3 Changed 7 years ago by benj

RINEX data downloaded from Thelma.

comment:4 Changed 7 years ago by ella

Applanix navigation processing now done. Files located in ARSF workspace.

comment:5 Changed 7 years ago by jb362

Started hyperspectral processing

comment:6 Changed 7 years ago by jb362

I need to check the Sbet file times against the flight line times in PosGNSS for this project as they don't seem to match the logsheet, however the extract directory on workspace is empty so there is no saved PosPac/PosGNSS data and no raw files. I think Loobos is going to have to be re-downloaded so that I can put the raw files back into workspace and then open up the project in PosGNSS

comment:7 Changed 7 years ago by jb362

downloading raw files done

comment:8 Changed 7 years ago by jb362

  • Owner set to jb362
  • Status changed from new to assigned

comment:9 Changed 7 years ago by jb362

The -1.nav didn't contain the navigation data but we have used the datacube.nav file and renamed it as -1.nav. Original -1.nav file renamed as -1.nav_old.

comment:10 Changed 7 years ago by jb362

.raw file missing for first eagle line

comment:11 Changed 7 years ago by jb362

No eagle data for flight line 1 or 9 and no hawk data for flight line 9 because of missing .raw files

comment:12 Changed 7 years ago by jb362

Actually .raw files do exits for flight line 9 on each sensor... I don't know why neither produced .tif files

comment:13 Changed 7 years ago by jb362

Offset times:

h1: 0.01
h2: 0.03
h3: 0.01
h4: 0.01
h5: 0.01
h6: 0.03
h7: 0.01
h8: 0.01
h9: Missing flightline
h10: 0.01
h11: 0.01
h12: 0.00
h13: 0.03
h14: -0.01
h15: 0.01
h16: 0.01
h17: 0.01
h18: 0.01
h19: 0.01
h20: 0.02
h21: 0.02
h22: 0.01
h23: 0.01
h24: 0.01
h25: -0.01
h26: 0.01
h27: 0.01
h28: 0.00
h29: 0.01

e1: Missing flightline
e2: 0.03
e3: 0.04
e4: 0.01
e5: 0.01
e6: 0.04
e7: 0.03
e8: 0.05
e9: Missing flightline
e10: 0.01
e11: 0.01
e12: 0.01
e13: -0.01
e14: -0.01
e15: 0.03
e16: 0.02
e17: 0.00
e18: 0.04
e19: 0.03
e20: 0.03
e21: 0.03
e22: 0.00
e23: 0.00
e24: 0.01
e25: 0.04
e26: -0.05
e27: 0.05
e28: 0.01
e29: 0.04

comment:14 Changed 7 years ago by jb362

Hyperspectral complete and waiting for delivery check

comment:15 Changed 7 years ago by chrfi

handing lidar over to knpa

roll offsets found so far

1202250.00093198
1211550.00090198
1342270.00093198
1347580.00093198
1337000.00096198
1326130.00093198
1221440.00093198
1351380.00093198
1358570.00093198
1404410.00093198
1410260.00093198
1416060.00093198
1421410.00090198
1433040.00093198

comment:16 Changed 7 years ago by knpa

  • Other processors set to knpa

Taking over responsibility for LIDAR.

comment:17 Changed 7 years ago by iopa

Performing delivery check on hyperspectral.

comment:18 Changed 7 years ago by iopa

Results:

  1. PI name spelling in Readme front page --> should be N. Kljun
  1. Flight line 9 for eagle and hawk has raw files and lev1's, so should produce .tif's as well.
  1. Since line 1 for eagle does not have a raw file, the rest of the lines will have to shift up so there is a continuous set of lines 1-28 for eagle. For the readme, this means that the lev1's and tif's 1-28 will need to be matched with the recorded line numbers in the logsheet for lines 2-29 (1 will be 004, 2 will be 019 etc.). This therefore applies to lev1's, lev3's and readme which will need to be renamed. (all this assuming eagle line 1 can not have a lev1 and lev3)
  1. Remove any hidden files in the delivery, currently there is just .Read_Me-20100722.tex.swp in EU10-01.
  1. Remove the .tex after the pdf is produced. Also remove the .tex in the screenshots directory.
  1. Missing bin directory, with instructions on downloading azgcorr (see http://arsf-dan.nerc.ac.uk/trac/wiki/Processing/FilenameConventions#Delivereddata)

comment:19 Changed 7 years ago by jb362

Made delivery check fixes

comment:20 Changed 7 years ago by jb362

Additional SCT offsets:

e9a: 0.01
e9b: 0.01
h9a: 1.30
h9b: 1.55

comment:21 Changed 7 years ago by jb362

  • Description modified (diff)

comment:22 Changed 7 years ago by jb362

If hawk data has not been rectified in two weeks time then eagle data should be delivered separately.

comment:23 Changed 7 years ago by jb362

Eagle SCT offsets:
e1: 0.04
e2: -0.01
e3: 0.08
e4: -0.10
e5: -0.05
e6: -0.08
e7: 0.09
e8: -0.18
e9: 0.28
e10: -0.30
e11: 0.09
e12: 0.05
e13: -0.06
e14: -0.10
e15: 1.26
e16: 0.00
e17: 0.03
e18: -0.07
e19: 0.03

The large SCT offset for flight line 15 is due to a sync error.

The header file for eagle flight line 16 had to be corrected so that "lines = 9596", rather than "lines = 1837" as in the original header file.

comment:24 Changed 7 years ago by jb362

Sorry, these offsets are for Santander (!!!)

comment:25 Changed 7 years ago by emca

Starting camera processing

comment:26 Changed 7 years ago by knpa

Line Roll
115710 0.00093198
120225 0.00093198
120723 0.00093198
121155 0.00090198
121650 0.00093198
122144 0.00093198
122708 0.00093198
123230 0.00093198
123725 0.00093198
124233 0.00093198
124752 0.00093198
125357 0.00093198
125940 0.00093198
130510 0.00093198
131013 0.00093198
131607 0.00093198
132108 0.00093198
132613 0.00093198
133147 0.00093198
133700 0.00096198
134227 0.00093198
134758 0.00093198
135318 0.00093198
135857 0.00093198
140441 0.00093198
141026 0.00093198
141606 0.00093198
142141 0.00090198
143304 0.00093198

comment:27 Changed 7 years ago by emca

Camera processing complete.
Removed photographs 1-11 as they did not cover the study area. Also removed 22, 44, 135, 136, 156, 202, 203, 221, 238, 239, 284, 285 & 292 due to under/over exposure.

Ready for delivery check at:
~airborne/workspace/EU10_01-2010_168_Loobos/delivery/20100824

comment:28 Changed 7 years ago by jb362

Delivery checking camera.

comment:29 Changed 7 years ago by jb362

Everything looks fine. Should PI on Readme just be N. Kljun or Dr N. Kljun?

comment:30 Changed 7 years ago by benj

Doesn't really matter. Pick one.

comment:31 Changed 7 years ago by emca

Photography data has been sent to PI 26/08/2010 along with LiDAR and Eagle data from day 168 (ticket #345)

comment:32 Changed 7 years ago by knpa

Finished processing LIDAR, creating delivery.

comment:33 Changed 7 years ago by jb362

Eagle data delivered 01/09/2010.

comment:34 Changed 7 years ago by ella

Starting LiDAR delivery check.

comment:35 Changed 7 years ago by ella

LiDAR check complete.

-Flight logsheet misspells site name.

-Read_me document figure 1 is intensity image without vectors but the caption says it has vectors overlaid.

comment:36 Changed 7 years ago by knpa

Fixed above issues, preparing delivery to PI.

comment:37 Changed 7 years ago by knpa

LIDAR delivered today (06/09/10) to N. Kljun at Swansea University.

comment:38 Changed 7 years ago by knpa

  • Description modified (diff)

comment:39 Changed 7 years ago by knpa

  • Description modified (diff)

comment:40 Changed 7 years ago by jb362

  • Description modified (diff)

comment:41 Changed 6 years ago by knpa

  • Description modified (diff)
  • Owner changed from jb362 to knpa

Has been transferred to repository and can be r-synced.

comment:42 Changed 6 years ago by emca

LAS files were re-processed using the above roll values and classified using TerraScan.
These .LAS files were then made available to PI via FTP 29/10/2010.

comment:43 Changed 6 years ago by knpa

Creating hawk delivery.

comment:44 Changed 6 years ago by knpa

Hawk reprocessing didn't include dem so may be some distortion in screenshots.

One of the lines didn't process to level three due to sync error. There is therefore a screenshot missing as well (line 9)

comment:45 Changed 6 years ago by knpa

SCT for line 9 given further up as:

h9a: 1.30
h9b: 1.55

Not sure what this means...will try a range between those two figures to determine which one fits best.

comment:46 Changed 6 years ago by knpa

Had to reprocess again once more due to incorrect bands given in raw header file (one of the 233 band ones).

As well as a sync error, h9 also fails because of a turn.
Benj says that the two scts above given for this line might be because it was split in two.

Finally, found a dem in delivery directory so don't need to worry about distorted screenshots any more.

comment:47 Changed 6 years ago by knpa

There are indeed different SCTs for the two parts that line 9 needs to be separated into.
However I found them to be:

h9a: 1.30
h9b: 1.74

It is strange that the values are different. Going ahead with delivery for now but this should be investigated at some point.

comment:48 Changed 6 years ago by knpa

Hawk ready to be delivery checked.

comment:49 Changed 6 years ago by emca

Delivery checked Hawk:

Readme:

  • Readme is named -old.pdf - should this be replaced with one in the main processing folder? This one is also for the Hawk data but the mosaic jpg is not correct (looks like its for a different project)
  • Readme says SRTM DEM is included but it is not in the delivery folder. If DEM is going to be included add it to the azgcorr command & delivery contents
  • Example azgcorr command section - remove comma in "using a pixel size of 2.0m, ."
  • Should mention over/under flows in readme

Logsheet needs to be renamed from 168-10_Loobos.pdf to EU10_01-2010_168_Loobos.pdf to conform with convention

Otherwise ready to delivery.

comment:50 Changed 6 years ago by knpa

Above fixes made.

comment:51 Changed 6 years ago by knpa

Recreating level 1's due to calibration mistake.

comment:52 Changed 6 years ago by knpa

Lev1's regenerated - awaiting checking.

comment:53 Changed 6 years ago by emca

Delivery checked re-processed hawk data. Data itself looks good. Ready to deliver once below items are addressed:

Updated readme needs to be moved from processing folder into delivery directory
Readme does not mention hawk bad pixel file
Readme quality section: entirity looks wrong, should it be entirety?
The PI and site name are incorrectly spelt on the logsheet
Bad pixel file is the old (incorrect) version - needs to be updated
Lev1 bil files need to be chmod'd to make them readable to all

comment:54 Changed 6 years ago by knpa

Above fixes made.

comment:55 Changed 6 years ago by emca

  • Description modified (diff)

Hawk data dispatched 21/12/2010

comment:56 Changed 5 years ago by mark1

When Eagle and Hawk are reprocessed in APL please notify PI and make available on ftp.

comment:57 Changed 5 years ago by chho

Hyperspectral delivery check: no problems, ready to deliver.

comment:58 Changed 5 years ago by adbe

Reprocessed hyperspectral delivered to Natascha Kljun on 10/09/2012 on disk 130

comment:59 Changed 5 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Milestone set to 2010 data processing completion

Beginning archiving (of original processing)

Note: See TracTickets for help on using tickets.